OpenStudy (anonymous):

Solve the system of equations and enter the solution as an ordered pair. 3x + 4y = 38 5x - 5y = -30

OpenStudy (anonymous):

5x-5y=-30 so, dividing both sides by 6: x-y=-6 x=y-6 Now plug x=y-6 into the first equation. 3(y-6) + 4y = 38 Solve for y. when you have y, you can find x using: x=y-6

OpenStudy (anonymous):

second eq by 5 x-y=-6 and into 4 4x-4y=-24 now add one and two 7x=14 so x=2 and from second 2-y=--5 y=7